Morro Bay SUP, LLC manufactures fiberglass paddleboards. The standard cost of direct materials and direct manufacturing labor is $100 per board. This includes 20 pounds of direct materials, at the budgeted price of $2 per pound, and five hours of direct manufacturing labor, at the budgeted rate of $12 per hour. Following are additional data for the month of March:

Units completed

6,000 units

Direct materials purchased

150,000 pounds

Cost of direct materials purchased

$292,500

Actual direct labor hour logged

32,000 hours

Actual direct labor cost

$368,000

Direct materials quantity variance

$12,500 U

Assuming there were no beginning inventories, calculate the following for March:

  1. DL variances
  2. Actual pounds of direct materials used in production
  3. Actual price per pound of direct materials purchased
  4. DM price variance
Answer #1

Direct labor variance = Standard labor cost-actual labor cost = (6000*5*12)-368000 = 8000 U

b) Direct material quantity variance = (Standard quantity-actual quantity)Standard price

-12500 = (6000*20*2-2X)

-12500 = 240000-2X

2X = 252500

X(actual pound) = 252500/2 = 126250 Pounds

c) Actual price per pound = 292500/150000 = 1.95 per pound

d) Direct material price variance = (Standard price-actual price)actual quantity = (2-1.95)*150000 = 7500 F